OpenStudy (anonymous):

Solve the system of equations: 4x = y - 1 6x - 2y = -3

hero (hero):

y = 4x+1 y = (3+6x)/2 y = y What happens when y = y?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

sry

hero (hero):

You set y = y and solve for x

hero (hero):

so: 2(4x+1) = 3 + 6x 8x + 2 = 3 + 6x 2x = 1 x = 1/2 y = 3

OpenStudy (anonymous):

oh ok i remember learning this. now i think i get it. Thank you so much hero

OpenStudy (anonymous):

X=.5 ; Y=3 (4x-y =-1 ) *2 -- eq 1 (8x-2y=-2)*-1 -- eq 2 solve 1&2

OpenStudy (anonymous):

but on this oneSolve the system of equations: 3x + y = -8 -2x - y = 6 i kept getting wrong though and i followed the steps

hero (hero):

you can do this one by combining them and eliminating y for the first step If you combine the equations you get x = -2, y = -2

hero (hero):